OpenStudy (anonymous):

Can someone tell me how you would simplify 16^1/2

OpenStudy (mathstudent55):

Here's an important rule: \(\Large{ a^{\frac{1}{n}} = \sqrt[n]{a}} \)

OpenStudy (anonymous):

That means nothing to me lol.

OpenStudy (anonymous):

@xandrineguillo , do you know this?

OpenStudy (mathstudent55):

Sometimes an example is helpful. \(9^{\frac{1}{2}} = \sqrt[2]{9} = 3\)

OpenStudy (anonymous):

@aliciabennet yes listen to mathstudent hes telling you the solution and the answer would be 4 its a matter of exponential laws

OpenStudy (anonymous):

Where did the 2 some from in step 2

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@xandrineguillo , your answer was WRONG

OpenStudy (mathstudent55):